OpenStudy (anonymous):

Multiplying Radical Expressions Simplify: the square root of 6 (7 the square root of 3 + 6) answer choices below 7 the square root of 108 21 the square root of 2 + 6 the square root of 6 7 the square root of 18 + 6 the square root of 6 126 the square root of 2

OpenStudy (campbell_st):

\[\sqrt{6}(7\sqrt{3} + 6)\] expanding gives \[7\times \sqrt{3\times6} + 6\times \sqrt{6}\] \[7\sqrt{18} + 6\sqrt{6}\] just simplify the 1st radical

OpenStudy (anonymous):

oh okay. thanks

OpenStudy (campbell_st):

the 2nd and 3 answers are the same... just written differently... the correct answer will depend on you are asked to completely simplify or not